Express as cm using decimals
60 mm
 Express as cm using decimals
60 mm

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

We know that there are 10 mm in 1 cm
60 mm = 60 / 10 cm
= 6.0 cm

Thanks
Shraddha Singh
3 years ago
Dear Student,
 
We know that,
 
10 mm = 1cm
 
1mm = 1/10 cm
 
60 mm = 60×1/10 cm
             = 6 cm
 
Hope it helps!!
Thank you ✌️